OpenStudy (anonymous):

freaking 2 equations in 2 unknowns with this problem x+y=4 x-y=2

OpenStudy (anonymous):

We know: \[ x+y=4\\ x-y=2 \]So: \[ x=2+y \]Substituting for \(x\) in #1: \[ (2+y)+y=4\implies\\ 2+2y=4\implies\\ 2y=2\implies\\ y=1 \]Plugging it back into our other equation: \[ x=2+(1)=3 \]Et voilá.

OpenStudy (anonymous):

thank you!